Week of 201102

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Kate (chair, DB), Borja (monitoring), Xavier (KIT), Vincent (security), Pavlo (ATLAS), Onno (NL-T1), Elena (CNAF), Vladimir (LHCb), David B (IN2P3), Andrew (TRIUMF), Martin (computing), Darren (RAL), Xin (BNL), Christoph (CMS), Dave M (FNAL), Pepe (PIC)

Experiments round table:

  • CMS reports ( raw view) -
    • Good CPU utilisation (CMS scales) 255k on average (~200k production, 55k analysis)
    • No major trouble

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC: mc
  • BNL: NTR
  • CNAF: we report attacks on T1 by Thai ip (202.29.220.182 and 202.29.214.13), we have requested the ban of those ip at NREN level
  • EGI: nc
  • FNAL: NTR
  • IN2P3: with current lockdown, IN2P3-CC is operated remotely since today for at least until 1st December. Nothing else to report.
  • JINR: NTR
  • KISTI: nc
  • KIT: We kindly ask to recalculate Alice's availability for last week (GGUS:149294). The ARC-CEs are still factored into their report, even though the service was officially decommissioned with our site downtime on October 6th. To prevent issues like that going forward we've also deleted the service endpoints from GOC-DB, which is why the decommission downtime for them no longer shows up.
  • NDGF: nc
  • NL-T1:
  • NRC-KI: nc
  • OSG: nc
  • PIC: NTR
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: NTR
  • CERN storage services: nc
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • Draft availability Sitemon reports for October sent around
    • XrootD restart campaign in good track, very little servers missing and restarts already agreed with managers
  • MW Officer: nc
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2020-11-11 - AndrewWong
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback